# AI-boosted rare event sampling to characterize extreme weather

### v1

@@ Discussion
To address the challenge posed by the scarcity of data for studying extreme events, we introduced AI+RES, a novel framework that synergistically combines RES, AI emulators, and GCMs such that they complement each other and alleviate each individual approach’s shortcomings. AI+RES solves the long-standing problem of finding an appropriate score function for RES in a high-dimensional chaotic system. As the first proof-of-concept, we applied this framework on mid-latitude heatwaves in the PlaSim GCM. We showed that this approach enables the sampling of extremely rare events and yields accurate estimates of their return periods up to 50,000 years with an ensemble size of only 400, with a reduction in computational cost of several hundred compared with direct sampling. We also showed that AI+RES significantly outperforms the current state-of-the-art RES that relies on simple score functions.
@@
Beyond the specific case of mid-latitude heatwaves, the potential applications of AI+RES are broad. It could easily be applied to other types of weather extremes—such as tropical cyclones, blocking events, precipitation, wind, or compound events. AI+RES offers a powerful tool to generate targeted catalogs of rare events in poorly sampled regions of the state space, particularly where direct sampling of computationally expensive GCMs offers insufficient coverage. The framework could also be well-suited to characterize the tails of distributions in sub-seasonal to seasonal ensemble forecasts, where large ensembles are needed. Moreover, by applying AI+RES to climate models running under different climate change scenarios (e.g., CMIP6 or the emerging global high-resolution runs), it would be possible to explore the evolution of the statistics of rare extremes in future climates. Note that the AI+RES can be readily applied to any state-of-the-art GCM; here, we focused the proof-of-concept on PlaSim as its computational efficiency allows generating a ground truth dataset.
@@ Acknowledgments
A.L. was funded in part by RTE France, the French transmission system operator, and benefited from the national agency for research and technology (ANRT) funding under the CIFRE contract 2023/0506. A.W. is supported by the Eric and Wendy Schmidt AI in Science Fellowship, a program of Schmidt Sciences. A.W. and A.L. are also supported by France-Chicago Center FACCTs Award 300004248. This work was also supported by the Institut des Mathématiques pour la Planète Terre (IMPT) and funded in part by the US National Science Foundation through award NSF RISE–2425898 and NSF-2531264. In addition, the authors thank Alessandro Lovo for fruitful discussions and insightful ideas on efficiently coupling RES and AI forecasts, which were invaluable to this project, Anna Asch for her work in making the diagnostic plots of the AI emulator, and Justin Finkel for engaging and insightful discussions. Computational resources were provided by NSF ACCESS (allocation ATM170020), NCAR’s CISL (allocation URIC0009), and the University of Chicago Research Computing Center.
@@
METHODS
GCM and Data
PlaSim [40] is an efficient general circulation model (GCM) of intermediate complexity that includes a spectral dynamical core that solves the primitive equations for vorticity, divergence, temperature, and humidity. We use PlaSim in this study because it allows rigorous validation of our methodology with large, ground-truth ensemble simulations and it has previously been employed in various studies as a benchmark model to explore the statistics of persistent heat extremes [18, 22], as well as for testing deep learning-based forecasting approaches for such extremes [44, 52]. PlaSim is coupled with simplified representations of land, ocean, and sea ice boundary layers. We use a T42 spectral truncation mapped onto a $64\times 128$ Gaussian grid, with 10 vertical levels. We run PlaSim with sea surface temperature and sea ice thickness fixed to a climatological, annually repeating cycle derived from the AMIP-II boundary dataset (1870–2006) [53], via linear interpolation of monthly climatologies. In the land model, surface temperature is computed through a linear energy balance approach, and soil hydrology is represented using a bucket model with regionally varying water holding capacity[54].
AI Weather Emulator
$\tilde{X}_{\text{atm},t+\Delta t}=\mathcal{F}_{\Theta}(X_{\text{atm},t},X_{\text{bnd},t}).$ (2)
where $\Theta$ are the trainable parameters of the neural network $\mathcal{F}_{\Theta}$. After initialization, the AI emulator may be cycled autoregressively to obtain long-term predictions. While the AI emulator is trained as a deterministic forecast model, we can generate ensemble forecasts by perturbing the initial conditions.
Architecture
We adopted the PanguWeather 3D Earth-specific transformer (EST) architecture for our AI emulator, which is specifically tailored for global geophysical fields by incorporating Earth-specific inductive biases such as spherical positional encoding and longitudinal periodicity. This architecture has demonstrated strong skill in medium-range forecasting as well as stability in long-term autoregressive simulations [55]. To adapt our emulator to the PlaSim GCM, we made a number of modifications to the original EST used in PanguWeather. To account for the decrease from PanguWeather’s $0.25^{o}$ horizontal resolution to Pangu-PlaSim’s $2.8^{o}$, we decreased the horizontal patch size from $(4,4)$ to $(2,2)$ and the horizontal attention window size from $(6,12)$ to $(2,4)$. We additionally increased the vertical attention window size to include all tokens to account for the inclusion of the top-of-atmosphere boundary variables. The original patch recovery layer was replaced with a pixel shuffle deconvolution layer, as in the ArchesWeather model of [47], to mitigate artifacts arising from patch recovery, while an additional convolutional layer for each variable type (3D prognostic, 2D prognostic, and 2D diagnostic) was added following patch recovery to allow for additional processing of the full-resolution recovered variables.
Training and Validation
Prior to input into the AI emulator, the atmospheric fields from the PlaSim GCM are vertically interpolated from the original 10 topography-following sigma levels to 13 equipressure surfaces. The details of these prognostic atmospheric variables, as well as the other variables input and output by the AI emulator, can be found in Extended Data Table 1. The AI emulator is trained to minimize a weighted sum of mean absolute errors between the model output and the prognostic and diagnostic variables at a timestep $\Delta t=6$ hours in the future. The precise weighted loss function is similar to that used in the PanguWeather model:
$\mathcal{L}(\tilde{X}_{\text{atm},t},X_{\text{atm},t})=\frac{1}{N_{a}N_{z}}\sum_{a,z}\left\|\tilde{X}_{a,z,t}-X_{a,z,t}\right\|_{1}+\frac{1}{4N_{s}}\sum_{s}\left\|\tilde{X}_{s,t}-X_{s,t}\right\|_{1}+\frac{1}{4N_{d}}\sum_{d}\left\|\tilde{X}_{d,t}-X_{d,t}\right\|_{1},$ (3)
To train the AI emulator, we use data from the final 100 years of a single 112-year-long PlaSim control run, and reserve year 11 for validation. Prior to input, this data is standardized using statistics from the 100-year training data set so that each variable at each pressure has a global mean of 0 and global standard deviation of 11. We additionally add a Gaussian noise vector independently sampled from $\mathcal{N}(0,2.5\times 10^{-3})$ to each standardized atmospheric variable input to the emulator during training. This noise addition mitigates an instability that can arise during autoregressive prediction with the AI emulator due to the reduced spectral resolution of the PlaSim GCM’s dynamical core relative to its spatial resolution. All other AI emulator training parameters and methods are taken from standard methods for training large-scale vision transformer-based models (see SI Table S1).
Ensemble Forecasting
Ensemble forecasts are generated by adding independent initial condition perturbations to all atmospheric variables at the beginning of the forecast. Perturbations are sampled from a Gaussian distribution with mean 0 and a standard deviation such that the resulting ensemble spread over time is similar to that of the PlaSim GCM with our selected magnitude of perturbations to the initial surface pressure (see SI Section S2).
Rare event Algorithm: Diffusion Monte Carlo
@@ METHODS
We use a version of the Diffusion Monte Carlo (DMC) algorithm that closely follows the formulations proposed in [19] and [20]. It relies on a one-dimensional score function (or reaction coordinate), $\theta:\mathbb{R}^{d}\to\mathbb{R}$, which assigns higher values to regions of phase space associated with the rare event of interest. The algorithm enhances sampling in regions where $\theta$ is high (via duplication or “splitting”) and suppresses sampling where it is low (via termination or “killing”).
@@
The mathematical properties of DMC have been rigorously studied—see, for instance, [56]—with results establishing its convergence and asymptotic behavior as the ensemble size NN tends to infinity. Under mild integrability assumptions, DMC provides unbiased estimators that converge in the large-NN limit. These convergence results are quite general and remain valid even for systems of arbitrarily high dimension $d$. Any statistical quantity that can be computed via direct sampling can, in principle, also be estimated using DMC. This includes observables that depend on the full trajectory of the process from the initial time $t_{0}$ up to a given time $t_{f}$. We denote by $\left(X_{t}^{n}\right)_{1\leq n\leq N}$ the NN realizations generated during the simulation and refer to each sample $X_{t}^{n}$ as a walker.
We introduce a finite sequence of scheduled resampling times $0=t_{0}<t_{1}<\cdots<t_{K}=t_{f}$, which may be spaced either uniformly or non-uniformly. For notational convenience, we set $X_{k}^{i}:=X_{t_{k}}^{i}$. The resampling steps are governed by a sequence of splitting functions $V_{k}$, which are themselves functions of the score function $\theta$.
@@
Algorithm 1 Diffusion Monte Carlo algorithm
Input: NN walkers $\left(X_{0}^{n}\right)_{1\leq n\leq N}$ starting from an initial condition.
@@
One should note that the DMC algorithm applies a selection procedure before the first simulation step. In practice, if one does not want to apply this first selection step, it can be avoided by choosing $V_{0}=0$. We do this in this study.
@@
where the expectation on the left is with respect to the distribution of the original process without RES. Using this formula with the choice $\psi=\mathbbm{1}_{A\geq a}$ for a large $a\in\mathbb{R}$ and an observable AA of the trajectory of $X$, one then has access to the probability of AA taking extreme values $\left(\mathbb{E}\left[\mathbbm{1}_{A\geq a}\right]=\mathbb{P}\left[A\geq a\right]\right)$. In our case, this observable is $A_{L}(t_{f})$ defined as the $L$-day average of the 2m temperature over the region $\mathcal{R}$ in Eq. (1).
Following [20], we choose a splitting function of the form $V_{k}=C_{k}\theta(X_{k})$, where $C_{k}>0$ is a constant that can vary at each resampling step. The choice of $C_{k}$ is critical as it controls the strength of the selection process; we discuss it in the Supplementary Information, Section S4.3.
To improve the algorithm’s robustness to the choice of $V_{k}$, [19] introduced an additional step known as quantile mapping, which gives the algorithm its full name: Quantile Diffusion Monte Carlo (QDMC). Further details can be found in the Supplementary Information, Section S4. In our work, however, we adopted a simpler alternative by rescaling the score function $\theta$ at each resampling step using the following transformation
$\tilde{\theta}_{k}(X^{i}_{k})=\frac{\theta(X^{i}_{k})-\mu_{\theta_{k}}}{\sigma_{\theta_{k}}},$ (5)
where $\mu_{\theta_{k}}$ and $\sigma_{\theta_{k}}$ denote the mean and standard deviation of $\theta$ across all walkers at time $t_{k}$. The rescaled score function $\tilde{\theta}_{k}$ is then used to compute the splitting function via $V_{k}=C_{k}\tilde{\theta}_{k}$. Using this approach, we obtained results that were very similar to those achieved with the quantile-mapping step.
Choice of the score function
At each resampling time and for each walker $i$, we perform an ensemble forecast from the AI emulator with $M$ members starting from the last state of the walker, until the target period $t_{f},t_{f}+L]$. This yields an ensemble of M forecasts ${\hat{A}^{i,j}_{L}(t_{f}|X^{i}_{k})}_{j=1}^{M}$ of the observable $A^{i}_{L}(t_{f})$ (see Eq. (1)) starting from the initial condition $X^{i}_{k}$.
@@
We note that $A^{i}_{L}(t_{f})$ could be predicted by other methods. We choose to leverage the AI emulator because of its proven forecasting skill and numerical efficiency, but other methods could be used. In particular, it would be interesting to test simpler methods based on directly learning the observable $A_{L}(t_{f})$ (or its distribution) from an initial condition, such as in [58, 59, 44, 60, 61, 62]. In this work, we also test the best possible forecast system, using ensembles of the PlaSim model itself to generate the score function as described in the Different baselines section.
@@
We note also that one could use a different statistic calculated from the ensemble forecast as the score function. The ideal score function to estimate the probability of achieving the rare event $A_{L}(t_{f})>a$ for large $a$ should anticipate walker paths leading to the rare event. An alternative score function appropriate for this goal is the conditional probability function $\theta(x)=P(A_{L}(t_{f})>a|X_{t_{k}}=x)$, which is referred to as the committor function [63]. This score function has been proven to be optimal for a splitting algorithm similar to quantile DMC [64]. However, in our case there is no clear threshold, $a$, since we aim to characterize the entire tail of the distribution of $A_{L}(t_{f})$. Our choice of score function, namely, the mean of the ensemble forecasts, favors the duplication of walkers for which the conditional expectation $\mathbb{E}[A_{L}(t_{f})\mid X_{t_{k}}=x]$ is large. This approach is more conservative than selecting a fixed threshold a priori.
Choice of other parameters
@@
The important parameters of the algorithm are given in Extended Data Table 2. We use $K=6$ steps with scaling constants $(C_{1},C_{2},C_{3},C_{4},C_{5},C_{6})\approx(0.,0.,0.,1.6,1.8,2.0)$, tuned empirically. The steps with $C_{k}=0$ ensure that the walkers are well-separated in phase space when selection begins.
Return period curves
$\mathbb{P}(A_{L}(t_{f})>a_{j})\approx\frac{1}{N}\sum_{i=1}^{N}\mathbbm{1}_{A_{L}(t_{f})>a_{j}}\left(Y_{t_{f}+T}^{i}\right)$ (7)
For the RES algorithm, we use the formula (4) with the observable $\psi(X)=\mathbbm{1}_{A_{L}(t_{f})>a_{j}}$ to compute the probability of exceedance. Namely,
$\mathbb{P}(A_{L}(t_{f})>a_{j})\approx\frac{\bar{w}_{K}}{N}\sum_{i=1}^{N}\mathbbm{1}_{A_{L}(t_{f})>a_{j}}\left(X_{t_{f}+T}^{i}\right)e^{-V_{K}\left(\hat{X}_{K}^{i}\right)}$ (8)
Extreme value theory fits
$F_{(\sigma,\xi)}(z)=1-\left[1+\xi\frac{(z-u)}{\sigma}\right]^{-1/\xi}_{+},$ (9)
To ensure a fair comparison in terms of sample size, the GPD distributions were fitted on datasets of the same size as the AI+RES experiments (namely $N=400$ in Fig. 2), on 100 independent training datasets. The gray shading represents the 10th–90th percentile range of return-period curves obtained from the GPD fits across these datasets.
In the Peak-Over-Threshold method, the key hyperparameter to tune is the threshold $u$ above which the data is modeled. We determine $u$ using the heuristic of the Mean Residual Life plot ([65]), selecting a percentile that can be applied consistently across all datasets. Based on this criterion, we retain the 90th percentile of the data as the threshold. Parameters were estimated using the scipy Python package ([66]).
Derivation of the computational speed-up factors
In the following, we derive two criteria to evaluate the computational gains provided by the AI+RES algorithm compared to DNS. We make the key assumption that the computational cost of running the AI emulator ensemble forecasts is negligible compared to that of running the physics-based model (in this case, the GCM). For reference, [7] report speed-ups of up to $\mathcal{O}(10^{4}\!-\!10^{5})$ and energy-consumption reductions of $\mathcal{O}(10^{4})$ for FourCastNet (AI emulator) relative to the IFS model (numerical weather prediction model). In our case, the cost comparison between the AI emulator and PlaSim is less favorable, as PlaSim is specifically designed to be computationally inexpensive. This is precisely what allows us to run a 50,000-member PlaSim ensemble within reasonable computational limits, enabling a robust validation of the methodology. Nevertheless, a preliminary analysis indicates that our emulator runs approximately 10 times faster on a single A100 GPU than PlaSim using 64 Intel Xeon Platinum 8160 CPU cores, even without any model variable or weight pruning, or inference-specific optimizations. This does not undermine our assumption of negligible cost for the AI emulator, as the present study serves as a proof of concept.
Sampling speed-up factor
$\text{SSUF}(a)=\frac{N_{DNS}(a)}{N_{RES}}=\frac{T_{a}n_{a}}{N_{RES}}$ (10)
We also introduced a metric to quantify the diversity of the sampled trajectories (Section S7 of the SI), demonstrating that the generated samples are not overly correlated.
Variance speed-up factor
The other main objective of RES is to reduce the uncertainty in rare probability estimates. Following [19, 20], we also quantify speed-up based on the reduction in the variance of estimated rare event probabilities. Suppose we have used RES to obtain an estimate of the probability of exceedance $\hat{p}_{a}$ above the value $a$ with a variance $\sigma_{RES}(a)^{2}$ using a budget of $N_{RES}$ walkers. We can compare this with the variance of the DNS estimator (Eq. (7)) as a function of NN and $a$, which is given by
$\sigma_{DNS}(a,N)^{2}=\frac{p_{a}(1-p_{a})}{N}$ (11)
$N_{DNS}(a)\approx\frac{\hat{p}_{a}(1-\hat{p}_{a})}{\hat{\sigma}_{RES}(a)^{2}}.$ (12)
We then define the Variance Speed-up Factor (VSUF) as
$\text{VSUF}(a)=\frac{N_{DNS}(a)}{N_{RES}}=\frac{p_{a}(1-p_{a})}{\hat{\sigma}_{RES}(a)^{2}N_{RES}}=\frac{T_{a}-1}{T_{a}^{2}\hat{\sigma}_{RES}(a)^{2}N_{RES}}$ (13)
In practice, we estimate the variance of RES, $\hat{\sigma}_{RES}(a)^{2}$, empirically from only 10 independent realizations of the algorithm, as running a larger number of experiments is computationally expensive. Consequently, the estimate of $\hat{\sigma}_{RES}(a)^{2}$ itself is subject to sampling error. In Section S8 of the SI, we also present alternative estimates of the variance using only a single realization of the algorithm.
@@
A large part of the applications of trajectory splitting-based RES algorithms in climate science has focused on sampling large deviations of long-time averages of an observable AA [18, 21, 22, 23], where the averaging window typically spans several months. In these approaches, the score function is usually constructed from instantaneous values of AA (or from its average over the most recent resampling interval) at the resampling time. While this strategy is effective for sampling long-time averages, it is not well-suited for sampling extreme events which occur on time scales shorter than the Lyapunov time of the system—about 5–10 days in the atmosphere. To address such events, a score function capable of anticipating extremes weeks in advance is required. This is precisely the approach taken in the present work, where forecasts from an AI-based emulator of the climate model are leveraged to construct the score function.
S2 How we generate PlaSim ensembles
As described in the Methods section, for the duplication of walkers at each resampling step in the DMC algorithm to have an effect, it is necessary to perturb their initial conditions. Following [18], we introduce perturbations by adding a small noise to the coefficients of the spherical harmonics of the logarithm of the surface pressure field. Fig. S1 shows the time evolution of the standard deviation of 100-member PlaSim ensembles generated with different values of $\epsilon$. In our DMC implementation, we use a perturbation amplitude of $\epsilon=3\times 10^{-3}$. We tested several values of $\epsilon$ and found that this choice provides a good balance: it enhances the ability of the PlaSim ensembles to explore the phase space, while remaining small enough to avoid altering the final statistics of the target observable or introducing non-physical discontinuities in the trajectories. We anticipate that more sophisticated perturbation strategies, such as Bred vectors [67] or early splitting [30], could further improve the performance of the algorithm.
S3 The Pangu-PlaSim Emulator
S3.1 Training hyperparameters

### v2

@@ Discussion
To address the challenge posed by the scarcity of data for studying extreme events, we introduced AI+RES, a novel framework that synergistically combines RES, AI emulators, and GCMs such that they complement each other and alleviate each individual approach’s shortcomings. AI+RES solves the long-standing problem of finding an appropriate score function for RES in a high-dimensional chaotic system. As the first proof-of-concept, we applied this framework on mid-latitude heatwaves in the PlaSim GCM. We showed that this approach enables the sampling of extremely rare events and yields accurate estimates of their return periods up to 50,000 years with an ensemble size of only 400, with a reduction in computational cost of several hundred compared with direct sampling. We also showed that, even with an imperfect emulator, AI+RES significantly outperforms the current state-of-the-art RES that relies on simple score functions.
@@
Beyond the specific case of mid-latitude heatwaves, the potential applications of AI+RES are broad. It could easily be applied to other types of weather extremes—such as tropical cyclones, blocking events, precipitation, wind, or compound events. AI+RES offers a powerful tool to generate targeted catalogs of rare events in poorly sampled regions of the state space, particularly where direct sampling of computationally expensive GCMs offers insufficient coverage. The framework could also be well-suited to characterize the tails of distributions in sub-seasonal to seasonal ensemble forecasts, where large ensembles are needed. Moreover, by applying AI+RES to climate models running under different climate change scenarios (e.g., CMIP6 or the emerging global high-resolution runs), it would be possible to explore the evolution of the statistics of rare extremes in future climates. Note that the AI+RES can be readily applied to any state-of-the-art GCM; here, we focused the proof-of-concept on PlaSim as its computational efficiency allows validation against a ground truth dataset.
Regarding the AI emulators themselves, the utility of AI+RES is twofold. As the climate modeling community moves toward building emulators for increasingly high-resolution GCMs, obtaining sufficient physically generated samples to accurately train and validate them with physically generated data becomes a key challenge for rare events. Using AI+RES to generate such catalogs of extremes will constitute precious training and validation datasets for the emulators.
@@ Acknowledgments
A.L. was funded in part by RTE France, the French transmission system operator, and benefited from the national agency for research and technology (ANRT) funding under the CIFRE contract 2023/0506. A.W. is supported by the Eric and Wendy Schmidt AI in Science Fellowship, a program of Schmidt Sciences. A.W. and A.L. are also supported by France-Chicago Center FACCTs Award 300004248. This work was also supported by the Institut des Mathématiques pour la Planète Terre (IMPT) and funded in part by the US National Science Foundation through award NSF RISE–2425898 and NSF-2531264. This research is also partially funded by the Institute for Climate and Sustainable Growth at the University of Chicago. In addition, the authors thank Alessandro Lovo for fruitful discussions and insightful ideas on efficiently coupling RES and AI forecasts, which were invaluable to this project, Anna Asch for her work in making the diagnostic plots of the AI emulator, and Justin Finkel for engaging and insightful discussions. Computational resources were provided by NSF ACCESS (allocation ATM170020), NCAR’s CISL (allocation URIC0009), and the University of Chicago Research Computing Center.
@@
End Matter
Definition of the heatwave index
Following previous studies [13, 45], we define the heatwave index $A_{L}(t)$ as a spatio-temporal average of 2-meter air temperature, T2m, as in the following:
$A_{L}(t):=\frac{1}{L}\int_{t}^{t+L}\left(\frac{1}{\mathcal{R}}\int_{\mathcal{R}}\text{T}_{2m}(\vec{r},s)\,d\vec{r}\right)ds.$ (1)
We want to estimate the return times of the events $A_{L}(t_{f})>a$ for large thresholds $a$. We typically use $L=7$ days, a window size small enough to capture the peak of mid-latitude heatwaves, yet large enough to have significant societal impacts. In this study, we consider two regions, $\mathcal{R}$, each defined as a $3\times 3$ grid point domain centered over France and Chicago (SM Fig. S7). We initialize simulations at $t_{0}=$ July 2 and integrate forward the PlaSim GCM until $t_{f}+L$, with $t_{f}=$ August 1, chosen to coincide with the climatological peak of $A_{L}(t_{f})$.
AI+RES with Diffusion Monte Carlo
@@
For the RES component of AI+RES, we use a version of the Diffusion Monte Carlo (DMC) algorithm that closely follows the formulations proposed in [14] and [16]. It relies on a one-dimensional score function (or reaction coordinate), $\theta:\mathbb{R}^{d}\to\mathbb{R}$, which assigns higher values to regions of phase space associated with the rare event of interest. The algorithm enhances sampling in regions where $\theta$ is high (via duplication or “splitting”) and suppresses sampling where it is low (via termination or “killing”).
@@
The mathematical properties of DMC have been rigorously studied — see, for instance, [53] — with results establishing its convergence and asymptotic behavior as the ensemble size $N$ tends to infinity. Under mild integrability assumptions, DMC provides unbiased estimators that converge in the large-$N$ limit. These convergence results are quite general and remain valid even for systems of arbitrarily high dimension $d$. Any statistical quantity that can be computed via direct sampling can, in principle, also be estimated using DMC. This includes observables that depend on the full trajectory of the process from the initial time $t_{0}$ up to a given time $t_{f}$. Although those convergence properties are very general, the choice of the score function can greatly influence the variances of the estimators obtained with DMC, highlighting the importance of this work.
We denote by $\left(X_{t}^{n}\right)_{1\leq n\leq N}$ the $N$ realizations generated during the simulation and refer to each sample $X_{t}^{n}$ as a walker. The DMC-based AI+RES splitting algorithm shown in Fig 1a proceeds as follows: we simulate $N$ walkers in parallel with the physics-based model (the PlaSim GCM, in our case), starting at time $t_{0}=0$ with small initial condition perturbations. An iterative duplication process is performed at fixed times $t_{1}<\cdots<t_{K}=t_{f}$, called resampling times and marked by dashed vertical lines in Fig. 1. These may be spaced either uniformly or non-uniformly; we chose here to perform resampling at uniform time intervals of length $\tau$, where $\tau$ is a tunable parameter. For notational convenience, we set $X_{k}^{i}:=X_{t_{k}}^{i}$. The resampling steps are governed by a sequence of splitting functions $V_{k}$, which are themselves functions of the score function $\theta$ (itself obtained from the AI emulator forecasts, in the case of AI+RES; see next section). Walkers with large scores are more likely to be duplicated, while walkers with small scores are more likely to be terminated. Following [16], we choose a splitting function of the form $V_{k}=C_{k}\theta(X_{k})$, where $C_{k}>0$ is a constant that can vary at each resampling step. In practice, the score function is rescaled before computing the splitting function, see SM Sec. S4.1 for more details. The choice of $C_{k}$ is critical, as it controls the strength of the selection process; we discuss it in SM Sec. S4.2. After resampling, we continue to simulate the selected walkers until the next resampling time, when the process is repeated, until the walkers reach the final time horizon $t_{f}=t_{K}$. Each walker is accompanied by a weight $w^{i}_{k}$ that allows unbiased estimation of rare event statistics of interest.
@@
Algorithm 1 Diffusion Monte Carlo (DMC) algorithm
Input: $N$ walkers $\left(X_{0}^{n}\right)_{1\leq n\leq N}$ starting from an initial condition.
@@
One should note that the DMC algorithm applies a selection procedure before the first simulation step. In practice, if one does not want to apply this first selection step, it can be avoided by choosing $V_{0}=0$. We do this in this work.
@@
where the expectation on the left is with respect to the distribution of the original process without RES. Using this formula with the choice $\psi=\mathbbm{1}_{A\geq a}$ for a large $a\in\mathbb{R}$ and an observable $A$ of the trajectory of $X$, one then has access to the probability of $A$ taking extreme values $\left(\mathbb{E}\left[\mathbbm{1}_{A\geq a}\right]=\mathbb{P}\left[A\geq a\right]\right)$. In our case, this observable is $A_{L}(t_{f})$ defined as the $L$-day average of the 2m temperature over the region $\mathcal{R}$ in Eq. (1).
Choice of the AI emulator-based score function
The AI emulator we use in AI+RES is trained to predict the state of the simulated physical system (the PlaSim GCM) a short time $\Delta t$ in the future given the state at time $t$, then cycled autoregressively for longer predictions. As shown in Fig. 1b, at each resampling time and for each walker $i$, we perform an $M$-member ensemble forecast using the AI emulator from the final state of the walker until the target period $t_{f},t_{f}+L]$. This yields an ensemble of $M$ forecasts ${\hat{A}^{i,j}_{L}(t_{f}|X^{i}_{k})}_{j=1}^{M}$ of the observable $A^{i}_{L}(t_{f})$ (see Eq. (1)) starting from the initial condition $X^{i}_{k}$. To the extent that the AI emulator faithfully reproduces the GCM’s dynamics, the prediction furnished by the AI emulator ensemble is ideally suited for use in constructing the score function. Our results show that the AI emulator’s weather forecast skill is sufficient to provide a very effective score function for AI+RES. More details on our AI emulator implementation are given in SM Sec. S3.
@@
We note that $A^{i}_{L}(t_{f})$ could be predicted by other methods. We choose to leverage the AI emulator because of its proven forecasting skill and numerical efficiency, but other methods could be used. In particular, it would be interesting to test simpler methods based on directly learning the observable $A_{L}(t_{f})$ (or its distribution) from an initial condition, such as in [55, 56, 45, 57, 58, 59]. As a baseline, we also test the best possible forecast system, using ensembles of the PlaSim model itself to generate the score function as described in the Ground truth and baselines section.
@@
We note that one could use a different statistic calculated from the ensemble forecast as the score function. The ideal score function to estimate the probability of achieving the rare event $A_{L}(t_{f})>a$ for large $a$ should anticipate walker paths leading to the rare event. An alternative score function appropriate for this goal is the conditional probability function $\theta(x)=P(A_{L}(t_{f})>a|X_{t_{k}}=x)$, which is referred to as the committor function [60]. This score function has been proven to be optimal for a splitting algorithm similar to quantile DMC [61]. However, in our case there is no clear threshold, $a$, since we aim to characterize the entire tail of the distribution of $A_{L}(t_{f})$. Our choice of score function, namely, the mean of the ensemble forecasts, favors the duplication of walkers for which the conditional expectation $\mathbb{E}[A_{L}(t_{f})\mid X_{t_{k}}=x]$ is large. This approach is more conservative than selecting a fixed threshold a priori.
.1 Choice of RES parameters
Params NN $M$ $K$ $\tau$ $C_{k}$ $L$
Values 400 100 6 5 days (0., 0., 0., 1.6, 1.8, 2.0) 7 days
@@
Table 1: Parameters used in the rare-event algorithm (Algorithm 1). NN is the number of walkers, $M$ the forecast ensemble size, $K$ the number of resampling steps, $\tau$ the resampling interval, $C_{k}$ the splitting constants, and $L$ the target duration. Simulations start on July 2 and run until $t_{f}+L$ with $t_{f}=$ August 1. The $C_{k}$ were tuned empirically, and steps with $C_{k}=0$ ensure that the walkers are well separated in phase space when selection begins.
@@
A large part of the applications of trajectory splitting-based RES algorithms in climate science have focused on sampling large deviations of long-time averages of an observable $A$ [13, 22, 12, 18], where the averaging window typically spans several months. In these approaches, the score function is usually constructed from instantaneous values of $A$ (or from its average over the most recent resampling interval) at the resampling time. While this strategy is effective for sampling long-time averages, it is not well-suited for sampling extreme events which occur on time scales shorter than the Lyapunov time of the system—about 5–10 days in the atmosphere. To address such events, a score function capable of anticipating extremes weeks in advance is required. This is precisely the approach taken in the present work, where forecasts from an AI-based emulator of the climate model are leveraged to construct the score function.
S2 PlaSim GCM
PlaSim [41] is an efficient general circulation model (GCM) of intermediate complexity that includes a spectral dynamical core that solves the primitive equations for vorticity, divergence, temperature, and humidity. We use PlaSim because it allows rigorous validation of our methodology with large, ground-truth ensemble simulations and it has previously been employed in various studies as a benchmark model to explore the statistics of persistent heat extremes [13, 12], as well as for testing deep learning-based forecasting approaches for such extremes [45, 62]. PlaSim is coupled with simplified representations of land, ocean, and sea ice boundary layers. We use a T42 spectral truncation mapped onto a $64\times 128$ Gaussian grid, with 10 vertical levels. We run PlaSim with sea surface temperature and sea ice thickness fixed to a climatological, annually repeating cycle derived from the AMIP-II boundary dataset (1870–2006) [63], via linear interpolation of monthly climatologies. In the land model, surface temperature is computed through a linear energy balance approach, and soil hydrology is represented using a bucket model with regionally varying water holding capacity[64].
S3 AI Weather Emulator
$\tilde{X}_{\text{atm},t+\Delta t}=\mathcal{F}_{\Theta}(X_{\text{atm},t},X_{\text{bnd},t}).$ (S1)
where $\Theta$ are the trainable parameters of the neural network $\mathcal{F}_{\Theta}$. After initialization, the AI emulator may be cycled autoregressively to obtain long-term predictions. While the AI emulator is trained as a deterministic forecast model, we can generate ensemble forecasts by perturbing the initial conditions. See Fig. S1 for an assessment of the forecast skill of the AI emulator.
S3.1 Architecture
We adopted the PanguWeather 3D Earth-specific transformer (EST) architecture for our AI emulator, which is specifically tailored for global geophysical fields by incorporating Earth-specific inductive biases such as spherical positional encoding and longitudinal periodicity. This architecture has demonstrated strong skill in medium-range forecasting as well as stability in long-term autoregressive simulations [65]. To adapt our emulator to the PlaSim GCM, we made a number of modifications to the original EST used in PanguWeather. To account for the decrease from PanguWeather’s $0.25^{\circ}$ horizontal resolution to Pangu-PlaSim’s $2.8^{\circ}$, we decreased the horizontal patch size from $(4,4)$ to $(2,2)$ and the horizontal attention window size from $(6,12)$ to $(2,4)$. We additionally increased the vertical attention window size to include all tokens to account for the inclusion of the top-of-atmosphere boundary variables. The original patch recovery layer was replaced with a pixel shuffle deconvolution layer, as in the ArchesWeather model of [48], to mitigate artifacts arising from patch recovery, while an additional convolutional layer for each variable type (3D prognostic, 2D prognostic, and 2D diagnostic) was added following patch recovery to allow for additional processing of the full-resolution recovered variables.
S3.2 Training and Validation
Prior to input into the AI emulator, the atmospheric fields from the PlaSim GCM are vertically interpolated from the original 10 topography-following sigma levels to 13 equipressure surfaces. The details of these prognostic atmospheric variables, as well as the other variables input and output by the AI emulator, can be found in Table S2. The AI emulator is trained to minimize a weighted sum of mean absolute errors between the model output and the prognostic and diagnostic variables at a timestep $\Delta t=6$ hours in the future. The precise weighted loss function is similar to that used in the PanguWeather model:
$\mathcal{L}(\tilde{X}_{\text{atm},t},X_{\text{atm},t})=\frac{1}{N_{a}N_{z}}\sum_{a,z}\left\|\tilde{X}_{a,z,t}-X_{a,z,t}\right\|_{1}+\frac{1}{4N_{s}}\sum_{s}\left\|\tilde{X}_{s,t}-X_{s,t}\right\|_{1}+\frac{1}{4N_{d}}\sum_{d}\left\|\tilde{X}_{d,t}-X_{d,t}\right\|_{1},$ (S2)
To train the AI emulator, we use data from the final 100 years of a single 112-year-long PlaSim control run, and reserve year 11 for validation. Prior to input, this data is standardized using statistics from the 100-year training data set so that each variable at each pressure has a global mean of 0 and global standard deviation of $1$. We additionally add a Gaussian noise vector independently sampled from $\mathcal{N}(0,2.5\times 10^{-3})$ to each standardized atmospheric variable input to the emulator during training. This noise addition mitigates an instability that can arise during autoregressive prediction with the AI emulator due to the reduced spectral resolution of the PlaSim GCM’s dynamical core relative to its spatial resolution. All other AI emulator training parameters and methods are taken from standard methods for training large-scale vision transformer-based models (see Table S1).
S3.3 Training hyperparameters